On January 1, 2018, the Highlands Company began construction on a new manufacturing facility for its own use. The building was completed in 2019. The company borrowed $1,500,000 at 10% on January 1 to help finance the construction. In addition to the construction loan, Highlands had the following debt outstanding throughout 2018:

$8,000,000, 15% bonds
$2,000,000, 10% long-term note


Construction expenditures incurred during 2018 were as follows:

January 1 $ 660,000
March 31 1,260,000
June 30 872,000
September 30 660,000
December 31 460,000


Required:
Calculate the amount of interest capitalized for 2018 using the specific interest method.

Date Expenditure Weight Average
January 1 x =
March 31 x =   
June 30 x =
September 30 x =
December 31 x =
Accumulated expenditure
Average Interest Rate Capitalized Interest
Average accumulated expenditures
x % =
x % =
Answer #1
Expenditure Weight Average
Jan-01 $             660,000 * 12/12 = $                      660,000
Mar-31 $          1,260,000 * 9/12 = $                      945,000
Jun-30 $             872,000 * 6/12 = $                      436,000
Sep-30 $             660,000 * 3/12 = $                      165,000
Dec-31 $             460,000 * 12/12 = 0
Accumulated expenditure $          3,912,000 $                  2,206,000
Average Interest rate Capitalised interest
Average accumulated expenditures $          2,206,000
Construction loan $          1,500,000 * 10% = $                      150,000
Other loans(not construction) $             706,000 * 14% = $                        98,840
$                      248,840
Working
Loan Amount Interest rate Amount
8000000 15% 1200000
2000000 10% 200000
10000000 1400000
Average rate 14.00%
1400000/10000000